### v1

Abstract
While the auto-bidding literature predominantly considers independent bidding, we investigate the coordination problem among multiple auto-bidders in online advertising platforms. Two motivating scenarios are: collaborative bidding among multiple distinct bidders managed by a third-party bidding agent, and strategic bid selection for multiple ad campaigns managed by a single advertiser. We formalize this coordination problem as a theoretical model and demonstrate that a straightforward coordination mechanism, where only the highest-value bidder competes with outside bids, strictly dominates independent bidding, improving both Return-on-Spend (RoS) compliance and the total value accrued for each participating auto-bidder or ad campaign. Additionally, our simulations on synthetic and real-world datasets support the theoretical result that coordinated mechanism outperforms independent bidding. These findings highlight both the theoretical potential and the practical robustness of coordination in auto-bidding in online auctions.
1 Introduction
Online advertising has been demonstrated as a prominent approach to monetize search engines, news feeds, social media, and e-commerce. Among them, real-time auctions have been used widely to connect advertisers and users efficiently. In the modern online advertising platforms, advertisers endeavor to maximize their campaign effectiveness, quantified by conversions or other pertinent metrics, through precise bid management that considers targeted return-on-spend (RoS) [aggarwal2019autobidding], which is also well-known as auto-bidding problems. To this end, a wide spectrum of bidding methodologies has been formulated, leveraging concepts from optimization theory, online learning paradigms, and game-theoretic principles [zhao2018deep, aggarwal2019autobidding, lee2013real, babaioff2021non, golrezaei2021bidding, deng2021towards, balseiro2021landscape, gao2022bidding, stram2024mystique].
Despite the wealth of sophisticated research in online advertising, existing literature predominantly concentrates on optimizing the value or utility from the perspective of individual bidders. This focus, while valuable, often overlooks a crucial aspect of real-world advertising dynamics: the potential for coordinated bidding strategies. In practice, it is common for multiple bidders to form coalitions, perhaps by authorizing a third-party advertising platform to manage their bids collectively. Alternatively, a single large advertiser, such as major e-commerce players like Amazon, Temu, or Shein, might control numerous distinct advertising campaigns. These entities often have a portfolio of advertisements they are willing to display to users, necessitating a coordinated approach across their various campaigns to maximize overall effectiveness, rather than optimizing each in isolation. Consequently, a deeper understanding of how auto-bidding algorithms can be coordinated across multiple allied bidders or campaigns, and what are the benefits of coordination over independent bidding, is becoming increasingly critical. Such understandings might be useful for industry practices and unlock further efficiencies in online advertising markets.
In this work, we investigate the coordination problem among multiple auto-bidding algorithms in repeated auctions. We formulate a theoretical model in which $N$ auto-bidders participate in $T$ rounds of second-price auctions. These bidders may form a coalition to determine their bids jointly, competing against bidders outside the coalition. Each auto-bidder is modeled as a value-maximizing agent subject to a Return on Spend (RoS) constraint (i.e., non-negative overall utility). We consider a simple form of coordination: only the bidder with the highest value inside the coalition submits a positive bid to compete against the outside bidders, while the other N1N-1 coalition members refrain from competing. In contrast, in the independent bidding scenario, all the $N$ bidders submit positive bids to compete against each other and the outside bidders. We aim to compare the coalition bidders’ welfare in the coordinated and independent bidding scenarios.
The main contributions of this paper are as follows. Theoretically, we prove that the highest-value-bidder-compete coordination mechanism mentioned above, although straightforward, enjoys a significant advantage over independent bidding: By coordination, the bidders in the coalition acquire higher values, as well as achieve lower RoS constraint violations, compared to independent bidding.
Specifically, we prove the following two results: (1) As long as the coalition bidders overbid (a common feature for value-maximizing bidding algorithms), then no matter what auto-bidding algorithms they use, coordination always increases the total utility (value - payment) of each bidder in the coalition over $T$ rounds, compared to the independent bidding scenario. This means that the bidders’ RoS constraints are better satisfied. (2) If coalition bidders run a state-of-the-art value-maximizing algorithm (a mirror-descent algorithm in feng2023online), then the total value obtained by the coalition is higher under coordination than under independent bidding.
Our theoretical results are proven under two mild assumptions. The first one (Assumption 1) is that the highest and second-highest values among the bidders inside the coalition should be competitive against the highest bid outside of the coalition. Without such an assumption, the coalition as a whole is too weak to compete against the outside bidders, and the problem becomes non-interesting. The second assumption is that the bidders in the coalition have independent and identically distributed (i.i.d.) valuations. This i.i.d. assumption is not only for theoretical simplicity but also reasonable in practice. For the two motivating scenarios mentioned above (multiple bidders forming a coalition and a single bidder controlling multiple advertising campaigns), the bidders/campaigns selected by the auctioneer to compete usually share similar features, justifying i.i.d. valuations. In addition, we show in Section 5 that our analysis can be extended to non-i.i.d. settings.
Finally, (in Section 6) we demonstrate by experiments, on both synthetic and real-world datasets, that our coordinated mechanism consistently outperforms independent bidding, validating our theoretical results under realistic market conditions.
1.1 Related Work
Bidder coordination, often referred to as cartels or collusion in economics, has been extensively studied in traditional auction theory [robinson1985collusion, hendricks1989collusion, marshall2007bidder]. Prior research has studied collusive behavior in first-price auctions [lopomo2011bidder, pesendorfer2000study], second-price auctions [mailath1991collusion, graham1987collusive], collusion-proof mechanism design [che2009optimal], auction design under collusion [pavlov2008auction], and collusion detection [chotibhongs2012analysis]. While these studies typically consider bidders forming collusions autonomously and reaching equilibrium outcomes in static auction settings [bergemann2016bayes, fu2025learning], we instead introduce a central planner that coordinates bidders to improve their total welfare in dynamic auction environments.
Recently, coordination behavior has also been observed in online advertising auctions [decarolis_bid_2023], sparking growing interest in understanding and analyzing such phenomena. decarolis2020marketing studied the impact of coordination on the revenue and efficiency of GSP and VCG auctions. romano2022power investigated the computational challenges faced by a media agency coordinating bidders under those mechanisms. chen_coordinated_2023 analyzed coordinated online bidding in repeated second-price auctions with budget constraints. While those works focus on coordinating utility maximizers, we study value maximizers. To the best of our knowledge, this is the first work to investigate coordination among value maximizers, which exhibit different properties from utility maximizers.
Our work is also related to the large literature on auto-bidding. For example, paes2024complex studied the dynamics of systems with multiple auto-bidders and revealed complex behaviors such as bi-stability, periodic orbits, and quasi-periodicity, while aggarwal2025multi analyzed optimal bidding strategies for advertisers across multiple platforms. For a comprehensive overview of the literature, we refer readers to the survey by aggarwal_auto-bidding_2024. Our study complements the predominant individual-optimization perspective of the auto-bidding literature.
2 Model: Coordinated Auto-Bidding
Repeated Second-Price Auctions with Coalition.
We consider the scenario where $N$ auto-bidders (or bidders) participate in $T$ rounds of repeated second-price auctions. In each round $t\in[T]=\{1,\ldots,T\}$, one item (e.g., ad slot) is for sale, and each bidder $i\in[N]=\{1,\ldots,N\}$ has a value $v_{i,t}\in[0,B]$ for the item i.i.d. sampled from a continuous distribution $F$. Let $\bm{v}_{t}=(v_{1,t},\ldots,v_{N,t})$ be the vector of values of all bidders at round $t$. Let $b_{i,t}$ denote bidder $i$’s bid at round $t$. The $N$ bidders may form a coalition. Our model captures two possible types of coalitions in ad auctions:
  1. 1.
    NN advertisers, each with an ad, instruct their auto-bidding algorithms to coordinate with each other before bidding.
  2. 2.
    One advertiser owns $N$ ads. Instead of using $N$ auto-bidding algorithms to bid for the $N$ ads independently, the advertiser instructs the algorithms to coordinate.
In either case, usually in practice, only similar advertisers/ads are selected to compete for one ad slot, which justifies the i.i.d. value assumption. In addition, we will show in Section 5 that our results extend to non-i.i.d. value settings.
There is a competing bid $d_{t}^{O}\in[0,B]$ from outside of the coalition, assumed to be i.i.d. sampled from another distribution $D$ every round. Our model also captures second-price auctions with reservation price, because $d_{t}^{O}$ can be the maximum between the reservation price and the outside competing bid. We denote the competing bid faced by bidder $i$ at round $t$ as
$d_{i,t}=\max\big\{d^{O}_{t},\max_{j\in[N]\setminus\{i\}}b_{j,t}\,\big\}.$ (1)
Let $x_{i,t}:=\mathbb{I}\{b_{i,t}\geq d_{i,t}\}\in\{0,1\}$ indicate whether bidder $i$ wins the item (ignoring tie-breaking for now), so bidder $i$’s payment is $p_{i,t}:=x_{i,t}d_{i,t}$ and utility is $u_{i,t}:=x_{i,t}(v_{i,t}-d_{i,t})$ at round $t$.
Value Maximization under RoS Constraint.
Each auto-bidder aims to maximize its total value while adhering to the Return on Spend (RoS) constraint. The formal definition of the optimization problem for bidder $i$ is as follows:
$\max_{\{b_{i,t}\}_{t=1}^{T}}$ $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}$ (bidder $i$’s value)
s.t. $\sum_{t=1}^{T}u_{i,t}=\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\sum_{t=1}^{T}p_{i,t}\geq 0$ (RoS constraint)
The RoS constraint ensures that for every dollar spent, at least one dollar of value is generated; it is equivalent to requiring the bidder’s total utility to be non-negative. [^1] In fact, the general RoS constraint requires $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\tau\cdot\sum_{t=1}^{T}p_{i,t}\geq 0$, i.e., at least $\tau$ dollar of value will be generated for every dollar spend, however, it is without loss of generality to assume $\tau=1$ in the above optimization problem because we can rescale the bidders’ values by dividing $\tau$.
Independent and Coordinated Auto-Bidding
Each auto-bidder runs an algorithm $A$ that learns to bid from history, in order to maximize the bidder’s total value subject to RoS constraint. Formally, at each round $t$, bidder $i$’s algorithm $A(\cdot)$ takes the historical information $H_{i,t}$ available to bidder $i$ (for example, the past values, bids, and allocations of the bidder: $H_{i,t}=\{v_{i,t^{\prime}},b_{i,t^{\prime}},x_{i,t^{\prime}}\}_{t^{\prime}=1}^{t-1}\cup\{v_{i,t}\}$) as input and outputs the bid $b_{i,t}$ for the current round. We refer to the scenario where the auto-bidding algorithms are run independently as independent bidding, formalized in Algorithm 1.
for $t=1,2,\dots,T$ do
   Each bidder $i\in[N]$ observes value $v_{i,t}$ and bids $b_{i,t}=A(H_{i,t})$;
   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$;
end for
ALGORITHM 1 Independent Bidding
Next, we define a coordinated bidding scenario where, at each round, only the highest-value bidder in the coalition places a positive bid while all other bidders bid zero. We assume that ties are resolved naturally. [^2] Ties happen with probability 0 given a continuous value distribution $F$. This is formalized in Algorithm 2.
for $t=1,2,\dots,T$ do
   Each bidder $i\in[N]$ observes value $v_{i,t}$;
   Let $i^{*}=\arg\max_{i\in[N]}v_{i,t}$;
   Bidder $i^{*}$ bids $b_{i^{*},t}=A(H_{i^{*},t})$; all $i\neq i^{*}$ bid $b_{i,t}=0$;
   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$;
end for
ALGORITHM 2 Coordinated Bidding
3 Coordination Reduces RoS Violation
In this section, we analyze the extent to which the RoS constraint of every auto-bidder is maintained/violated. We will show that, under mild assumptions, coordinated bidding results in a smaller RoS violation compared to independent bidding. Equivalently, every bidder in the coalition obtains a higher utility through coordination.
Formally, let $U_{i}^{C,A},U_{i}^{I,A}$ be the total utility of bidder $i$ during the $T$ rounds using auto-bidding algorithm $A$ under coordinated and independent bidding scenarios, respectively:
$U_{i}^{C,A}=\sum_{t=1}^{T}u_{i,t}^{C,A},\quad U_{i}^{I,A}=\sum_{t=1}^{T}u_{i,t}^{I,A}.$ (2)
RoS constraint requires $U_{i}^{C,A}\geq 0$ and $U_{i}^{I,A}\geq 0$. If $U_{i}^{C,A}\geq-c$ or $U_{i}^{I,A}\geq-c$, we say the RoS constraint for bidder $i$ is violated by $c$.
We allow the bidders to use any auto-bidding algorithm that (weakly) overbids: namely, $A(\cdot)$ can be an arbitrary function that satisfies $b_{i,t}=A(H_{i,t})\geq v_{i,t}$. Note that we are considering value-maximizing bidders in second-price auctions. If the bidders underbid ($b_{i,t}<v_{i,t}$), they will always obtain non-negative utility and satisfy the RoS constraint, but not necessarily maximize their total values. So we assume the bidders to overbid.
We make a reasonable assumption on the distributions of value and outside bid, $F$ and $D$. Let $(x)_{+}=\max\{x,0\}$ denote the positive part of a number. Let $v_{(N)},v_{(N-1)}$ be the largest and second-largest values among $N$ i.i.d. samples from $F$. We assume that: in expectation, the advantage of the second-largest value $v_{(N-1)}$ over the outside bid $d^{O}$ is larger than the advantage of $d^{O}$ over the largest value $v_{(N)}$:
Assumption 1
$\mathbb{E}_{F,D}\big[\big(v_{(N-1)}-d^{O}\big)_{+}-\big(d^{O}-v_{(N)}\big)_{+}\big]\geq\Delta>0$.
Intuitively, Assumption 1 says that the largest two values among the coalition bidders should be competitive enough against the outside bid. With large $v_{(N-1)}$ and $v_{(N)}$, $\mathbb{E}[(v_{(N-1)}-d^{O})_{+}]$ is large while $\mathbb{E}[(d^{O}-v_{(N)})_{+}]$ is small, which satisfy the assumption.
Example 1
The following examples satisfy Assumption 1:
  • $N=4,F=U[0,1],D=U[0,1]$, with $\Delta=1/6$.
  • $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$, with $\Delta=1/40$.
We will perform simulations on these examples in Section 6.
The main result of this section is the following theorem, which shows that coordinated bidding strictly improves each bidder’s total utility compared to independent bidding. This directly implies that each bidder’s RoS constraint violation under coordination is smaller.
Theorem 3.1(Coordination reduces RoS violation)
Assume Assumption 1. For any auto-bidding algorithm $A$ that overbids, for every bidder $i\in[N]$, with probability at least $1-\exp(-\frac{\Delta^{2}T}{32B^{2}N^{2}})$,
$U_{i}^{C,A}~\geq~U_{i}^{I,A}+\tfrac{\Delta T}{2N}.$
The proof of Theorem 3.1 involves an interesting comparison between the coordinated bidding scenario and the truthful bidding scenario where all bidders bid their values exactly, and a comparison between truthful bidding and independent auto-bidding. These two comparisons together imply that coordinated bidding generates higher utility for every coalition member than independent bidding. We present the detailed proof below.
3.1 Proof of Theorem 3.1
Let $U_{i}^{\mathrm{Truth}}=\sum_{t=1}^{T}u_{i,t}^{\mathrm{Truth}}$ be the total utility of bidder $i$ when all bidders in the coalition bid their values truthfully: $b^{\mathrm{Truth}}_{i,t}=v_{i,t}$. We will prove Theorem 3.1 by proving $U_{i}^{C,A}\geq U_{i}^{\mathrm{Truth}}+\frac{\Delta T}{2N}$ and $U_{i}^{\mathrm{Truth}}\geq U_{i}^{I,A}$. The second inequality is formalized below:
Lemma 1
For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, $U_{i}^{\mathrm{Truth}}\geq U_{i}^{I,A}$.
Proof
This lemma directly follows from the dominant-strategy-incentive-compatibility of second-price auction, which ensures that each bidder’s truthful-bidding utility $u^{\mathrm{Truth}}_{i,t}$ is weakly larger than the non-truthful bidding utility $u^{I,A}_{i,t}$ at every round. ∎
We then prove $U_{i}^{C,A}\geq U_{i}^{\mathrm{Truth}}+\frac{\Delta T}{2N}$:
Lemma 2
[…]
Proof
For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, consider the difference between bidder $i$’s coordinated auto-bidding utility and truthful bidding utility at round $t$:
$\Delta_{i,t}$ $=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$
$=(v_{i,t}-d_{t}^{O})\cdot\mathbb{I}[b_{i,t}^{C,A}\geq d_{t}^{O}]\cdot\mathbb{I}[v_{i,t}=\max_{j\in[N]}v_{j,t}]$
$\quad-(v_{i,t}-d_{i,t}^{\mathrm{Truth}})\cdot\mathbb{I}[v_{i,t}\geq d_{i,t}^{\mathrm{Truth}}]$
where $d_{i,T}^{\mathrm{Truth}}=\max\{d_{t}^{O},\max_{j\neq i}v_{j,t}\}$ is the competing bid for bidder $i$ under truthful bidding. We do a case analysis:
  • If $v_{i,t}\neq\max_{j\in[N]}v_{j,t}$, then bidder $i$ loses under both coordinated bidding and truthful bidding, so $u_{i,t}^{C,A}=u_{i,t}^{\mathrm{Truth}}=0$ and $\Delta_{i,t}=0$.
  • If $v_{i,t}=\max_{j\in[N]}v_{j,t}$, then
    $u_{i,t}^{C,A}=(v_{i,t}-d_{t}^{O})\cdot\mathbb{I}[b_{i,t}^{C,A}\geq d_{t}^{O}]$
    $u_{i,t}^{\mathrm{Truth}}=(v_{i,t}-d_{t}^{\mathrm{Truth}})\cdot\mathbb{I}[v_{i,t}\geq d_{t}^{O}].$
    Because the auto-bidder is overbidding ($v_{i,t}\leq b_{i,t}^{C,A}$), there are three further cases:
    • $d_{t}^{O}<v_{i,t}\leq b_{i,t}^{C,A}$: in this case, $\Delta_{i,t}=d_{t}^{\mathrm{Truth}}-d_{t}^{O}\geq 0$.
    • $v_{i,t}\leq b_{i,t}^{C,A}<d_{t}^{O}$: in this case, the bidder loses under both scenarios, so $\Delta_{i,t}=0$.
    • $v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}$: in this case, the bidder loses under truthful bidding but wins under coordinated bidding, so $\Delta_{i,t}=u_{i,t}^{C,A}=v_{i,t}-d_{t}^{O}\leq 0$.
Then, we condition on any history $H_{t-1}$, randomize over $\bm{v}_{t}$ and $d_{t}^{O}$, and analyze the expectation of $\Delta_{i,t}$. Let $F_{(N)}$ be the distribution of the largest value $v_{(N)}$ among $N$ samples from $F$.
[…]
$=\frac{1}{N}\bigg(\Pr_{v_{i,t}\sim F_{(N)},d_{t}^{O}\sim D}\Big[v_{i,t}>d_{t}^{O}\Big]\cdot\mathbb{E}\Big[d_{t}^{\mathrm{Truth}}-d_{t}^{O}\,\big|\,v_{i,t}\geq d_{t}^{O}\Big]$
$\quad+\Pr_{v_{i,t}\sim F_{(N)},d_{t}^{O}\sim D}\Big[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\Big]\cdot\mathbb{E}\Big[v_{i,t}-d_{t}^{O}\,\big|\,v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\Big]\bigg)$
$=\frac{1}{N}\bigg(\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}>d_{t}^{O}\big]\cdot\big(\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\big)\Big]$
$\qquad+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg).$
Because $\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}>0$ implies $\mathbb{I}\big[v_{(N)}>d_{t}^{O}\big]=1$, the above is equal to
$=\frac{1}{N}\bigg(\mathbb{E}\Big[\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\Big]$
$\qquad+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\leq b_{i,t}^{C,A}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg)$
$\geq\frac{1}{N}\bigg(\mathbb{E}\Big[\max\{v_{(N-1)},d_{t}^{O}\}-d_{t}^{O}\Big]+\mathbb{E}\Big[\mathbb{I}\big[v_{(N)}\leq d_{t}^{O}\big]\cdot\big(v_{(N)}-d_{t}^{O}\big)\Big]\bigg)$
$=\frac{1}{N}\bigg(\mathbb{E}\Big[\big(v_{(N-1)}-d_{t}^{O}\big)_{+}\Big]-\mathbb{E}\Big[\big(d_{t}^{O}-v_{(N)}\big)_{+}\Big]\bigg).$
Using Assumption 1, we obtain
[…]
Because values and outside bids are bounded by $B$, $\Delta_{i,t}$ is bounded by [B,B][-B,B]. So, by Azuma’s inequality,
[…]
Letting ${\varepsilon}=\frac{\Delta T}{2N}$ proves the lemma. ∎
Lemmas 1 and 2 together prove Theorem 3.1.
4 Coordination Increases Bidders’ Value
In this section, we show that coordination can increase the total value of auto-bidders, compared to independent bidding.
We will focus on a specific auto-bidding algorithm: the mirror-descent algorithm in feng2023online. It is the state-of-the-art algorithm for value maximization under RoS constraint, with near-optimal performances for the bidders, achieving $O(\sqrt{T})$ regrets. The algorithm maintains a parameter $\lambda_{i,t}$ for each bidder (which corresponds to the Lagrange multiplier of the dual optimization problem), so the bidder overbids by $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$ at each round. After observing a utility feedback, the bidder adjusts the parameter $\lambda_{i,t}$ in the opposite direction of the utility (i.e., positive utility decreases $\lambda_{i,t}$). The full algorithm is given in Algorithm 3.
Initialize multiplier $\lambda_{i,1}=1$, learning rate $\alpha=1/\sqrt{T}$;
for $t=1,2,\dots,T$ do
   Observe value $v_{i,t}$ and set bid $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$;
   Observe allocation $x_{i,t}$ and payment $p_{i,t}$, compute utility $g_{i,t}=v_{i,t}\cdot x_{i,t}-p_{i,t}$;
   Update the multiplier: $\lambda_{i,t+1}=\lambda_{i,t}\exp(-\alpha g_{i,t})$;
end for
ALGORITHM 3 Auto-bidding algorithm in feng2023online
For Algorithm 3, we show that coordinated auto-bidders obtain higher total value than independent auto-bidders. Let
$V_{t}^{C}=\sum_{i=1}^{N}v_{i,t}x^{C}_{i,t},\quad V_{t}^{I}=\sum_{i=1}^{N}v_{i,t}x^{I}_{i,t}$
be the total values obtained by the $N$ bidders at round $t$ in the coordinated case and independent case, respectively.
Theorem 4.1(Coordination Increases Total Value)
Assume 1, and suppose bidders run Algorithm 3. As $T\to\infty$, in expectation, the bidders’ total value under coordinated bidding is larger than that under independent bidding:
$\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{C}_{t}-\frac{1}{T}\sum_{t=1}^{T}V^{I}_{t}\Big]\geq 0.$
To prove Theorem 4.1, we first reduce the $N$-bidder coordinated scenario to a single-bidder problem where the single bidder’s value is the maximum among the coalition members. Using this reduction, we prove an interesting fact: under Assumption 1, the multipliers $\lambda_{i,t}^{C}$ of coordinated bidders converge to 0 in the limit as $t\to\infty$, meaning that the coordinated bidders will eventually submit arbitrarily high bids. This implies that the coordinated coalition will obtain a higher value than the independent case. We present the full proof below.
4.1 Proof of Theorem 4.1
Reduction to a single-bidder scenario.
To prove Theorem 4.1, we first draw a connection between the $N$-bidder coordination scenario and the scenario where a single bidder, whose value equals the highest value among the $N$ bidders, competes against the outside bid.
Formally, let $v_{(N)}=\max\{v_{1},\ldots,v_{N}\}$ denote the highest value among $N$ i.i.d. samples from $F$. Let $F_{(N)}$ be the distribution of $v_{(N)}$. Let $G(\lambda)$ be the expected utility of a single bidder with value $v_{(N)}\sim F_{(N)}$ and bidding using multiplier $\lambda>0$, competing against the outside bid $d^{O}\sim D$:
$G(\lambda)=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Big[(v_{(N)}-d^{O})\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big]\Big].$
Let $G_{i}^{C}(\lambda_{i})$ be the expected utility of bidder $i$ with value $v_{i}\sim F$ and multiplier $\lambda_{i}>0$ in the $N$-bidder coordinated scenario:
$G_{i}^{C}(\lambda_{i})=\mathbb{E}_{v_{1},\ldots,v_{N}\sim F,d^{O}\sim D}\big[(v_{i}-d^{O})x_{i}(\bm{v},\lambda_{i})\big]$
where $x_{i}(\bm{v},\lambda_{i})=1$ if $v_{i}=\max_{j\in[N]}\{v_{j}\}$ and $(1+1/\lambda_{i})v_{i}>d^{O}$, and 0 otherwise.
We show that the $N$-bidder coordinated scenario is “equivalent” to the single-bidder scenario in the following sense:
Observation 1
$G_{i}^{C}(\lambda_{i})=\frac{1}{N}G(\lambda_{i})$.
Proof
In the coordinated scenario, bidder $i$ obtains utility 0 if its value is not the highest among the $N$ bidders. The probability that bidder $i$ has the highest value is $1/N$ because the $N$ bidders’ values are i.i.d. So,
$G_{i}^{C}(\lambda_{i})$ $=0+\tfrac{1}{N}\mathbb{E}_{v_{i}\sim F|v_{i}=v_{(N)}}\big[(v_{i}-d^{O})\mathbb{I}[(1+1/\lambda_{i})v_{i}>d^{O}]\big]$
$=\tfrac{1}{N}\mathbb{E}_{v_{(N)}\sim F_{(N)}}\big[(v_{(N)}-d^{O})\mathbb{I}[(1+1/\lambda_{i})v_{(N)}>d^{O}]\big]$
$=\tfrac{1}{N}G(\lambda_{i}).$
Since the $N$-bidder coordinated scenario is equivalent to a single-bidder scenario, we can now analyze the single-bidder scenario. We present two useful lemmas regarding the expected utility function $G(\lambda)$ of the single bidder.
Lemma 3
$G(\lambda)$ is monotonically increasing in $\lambda$.
Proof
See Appendix 0.A. ∎
Lemma 4
Under Assumption 1, for any $\lambda>0$, $G(\lambda)\geq\Delta>0$.
Proof
By Lemma 3, $G(\lambda)$ is increasing in $\lambda$. So,
$G(\lambda)$ $\geq\lim_{\mu\to 0^{+}}G(\mu)$
$=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\big[v_{(N)}-d^{O}\big]$
$=\mathbb{E}\big[(v_{(N)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]$
$\geq\mathbb{E}\big[(v_{(N-1)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]\geq\Delta>0$
where the last $\geq$” follows from Assumption 1. ∎
Intuitively, Lemma 4 shows that the expected utility obtained by the highest-value coalition bidder is always positive (by a margin of $\Delta>0$). This means that the bidder’s multiplier $\lambda_{t}$, following mirror descent updates, will decrease in expectation every round and eventually converge to 0. We formalize this intuition below.
Coordinated multipliers converge to 0.
Using the above characterizations, we show an interesting fact: the multipliers $\lambda_{i,t}^{C}$ of auto-bidders running Algorithm 3 in the coordinated scenario converge to 0 as $t\to\infty$. This means that the coalition will eventually submit arbitrarily high bids $b_{i,t}^{C}=(1+\frac{1}{\lambda_{i,t}^{C}})v_{i,t}\to\infty$.
Lemma 5
Under Assumption 1, for any $i\in[N],t\in[T]$, with probability at least $1-\exp\big(-\frac{\Delta^{2}t}{32B^{2}N^{2}}\big)$, $\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\big(-\frac{\Delta t}{2N\sqrt{T}}\big)$.
Proof
According to Algorithm 3, $\lambda_{i,t+1}^{C}$ is equal to
$\lambda_{i,t+1}^{C}=\lambda_{i,1}\exp\Big(-\alpha\sum_{t^{\prime}=1}^{t}g_{i,t^{\prime}}\Big)$
where each $g_{i,t^{\prime}}$ is an unbiased estimator of bidder $i$’s expected utility $G_{i}^{C}(\lambda_{i,t^{\prime}})=\mathbb{E}[g_{i,t^{\prime}}]$ at round $t^{\prime}$. If bidder $i$ is not the highest-value bidder in round $t^{\prime}$, then $g_{i,t^{\prime}}=0$, meaning that the bidder’s multiplier $\lambda_{i,t^{\prime}}^{C}$ is not updated in that round. According to Observation 1 and Lemma 4, bidder $i$’s expected utility is at least
$\mathbb{E}[g_{i,t^{\prime}}]=G_{i}^{C}(\lambda_{i,t^{\prime}})=\frac{1}{N}G(\lambda_{i,t^{\prime}})\geq\frac{\Delta}{N}.$
Because values and outside bids are bounded by $B$, the realized utility $g_{i,t^{\prime}}\in[-B,B]$. So, by Azuma’s inequality:
$\Pr\left[\sum_{t^{\prime}=1}^{t}g_{i,t^{\prime}}<t\cdot\frac{\Delta}{N}-{\varepsilon}\right]\leq\exp\Big(-\frac{{\varepsilon}^{2}}{8B^{2}t}\Big).$
Let ${\varepsilon}=\frac{t\Delta}{2N}$. We have with probability at least $1-\exp(-\frac{t\Delta^{2}}{32B^{2}N^{2}})$,
$\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\Big(-\alpha\big(t\cdot\frac{\Delta}{N}-{\varepsilon}\big)\Big)=\lambda_{i,1}\exp\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big).$
where we plugged in $\alpha=1/\sqrt{T}$. ∎
Coordinated total value is larger than independent total value.
Using the fact that coordinated multipliers converge to 0, we show that the total value obtained by coordinated auto-bidders is larger than that under independent bidding in the limit $T\to\infty$.
Let $V_{(N)}(\lambda)$ be the expected value obtained by a single bidder whose value is the highest value $v_{(N)}$ among $N$ samples from $F$ and who uses multiplier $\lambda>0$ to bid against outside bid $d^{O}\sim D$:
$V_{(N)}(\lambda)=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Big[v_{(N)}\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big]\Big].$
Because $V_{(N)}(\lambda)$ is monotone in $\lambda$ and bounded, the limit
$V_{(N)}(0^{+})=\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$
exists. Note that the expected total value of $N$ coordinated bidders with multipliers $\lambda_{1}^{C},\ldots,\lambda_{N}^{C}$ is
$V^{C}(\lambda_{1}^{C},\ldots,\lambda_{N}^{C})=\mathbb{E}\Big[v_{i^{*}}\!\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda_{i^{*}}})v_{i^{*}}>d^{O}\big]\Big]=\mathbb{E}\big[V_{(N)}(\lambda_{i^{*}}^{C})\big]$
where $i^{*}=\operatorname*{arg\,max}_{i\in[N]}v_{i}$. This means that the expected average total value of the coordinated bidders is
$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{C}\Big]=\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i^{*}_{t},t}^{C})\Big]$ (3)
where $i^{*}_{t}\in\{1,\dots,N\}$ is the highest-value bidder in round $t$.
Let $V^{I}(\lambda_{1},\ldots,\lambda_{N})$ be the expected total value of $N$ independent bidders with multipliers $\lambda_{1},\ldots,\lambda_{N}>0$:
$V^{I}(\lambda_{1},\ldots,\lambda_{N})=\mathbb{E}_{\bm{v}\sim F,d^{O}\sim D}\Big[\sum_{i=1}^{N}v_{i}x_{i}(\bm{b}^{I})\Big]$
where $\bm{b}^{I}=\big((1+\tfrac{1}{\lambda_{1}})v_{1},\ldots,(1+\tfrac{1}{\lambda_{N}})v_{N},d^{O}\big)$.
Lemma 6
$V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$, for any $\lambda_{1},\ldots,\lambda_{N}>0$.
Proof
Pick $\lambda_{\varepsilon}<\min\{\lambda_{1},\dots,\lambda_{N}\}$. If some bidder $i\in[N]$ wins the item under independent bidding, then their bid $b_{i}^{I}=(1+\frac{1}{\lambda_{i}})v_{i}$ is larger than the outside bid $d^{O}$. In that case, suppose the highest-value bidder uses multiplier $\lambda_{\varepsilon}$, then their bid is $(1+\frac{1}{\lambda_{\varepsilon}})v_{(N)}\geq(1+\frac{1}{\lambda_{i}})v_{i}=b_{i}^{I}>d^{O}$, so the highest-value bidder also wins in the single-bidder scenario. Note that the value obtained by the highest-value bidder is at least $v_{i}$, so the expected value $V_{(N)}(\lambda_{\varepsilon})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$. Letting $\lambda_{\varepsilon}\to 0$ proves the lemma. ∎
Proof(Proof of Theorem 4.1)
By Lemma 5, with constant $c\coloneqq\frac{\Delta^{2}}{32B^{2}N^{2}}$, for any fixed $i$ and $t\geq 1$, the coordinated multiplier $\lambda_{i,t+1}^{C}$ satisfies
$\Pr\!\left[\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\!\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big)\right]\;\geq\;1-e^{-ct}.$ (4)
Fix $\varepsilon>0$. Since $V_{(N)}(0^{+})$ exists, there is $\delta=\delta(\varepsilon)>0$ such that
$\big|V_{(N)}(0^{+})-V_{(N)}(\lambda)\big|<\varepsilon\quad\text{for all }\lambda\in(0,\delta].$ (5)
Define
$m_{T}\;\coloneqq\;\left\lceil\frac{2N\sqrt{T}}{\Delta}\,\log\!\Big(\frac{\lambda_{i,1}}{\delta}\Big)\right\rceil~=~O\Big(\sqrt{T}\log\frac{1}{\delta}\Big),$
so for all $t\geq m_{T}$,
$\lambda_{i,1}\exp\!\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big)\leq\lambda_{i,1}\exp\!\Big(-\frac{m_{T}\Delta}{2N\sqrt{T}}\Big)\leq\delta.$
Let $E_{i,t}$ be the event inside the probability in (4). A union bound over $i=1,\dots,N$ and $t\geq m_{T}$ gives
$\Pr\!\Big[\bigcap_{i=1}^{N}\bigcap_{t=m_{T}}^{T}E_{i,t}\Big]\;\geq\;1-NTe^{-cm_{T}}\;\eqqcolon\;1-p_{T}.$ (6)
On this event, for all $t\geq m_{T}$ we have $\lambda^{C}_{i^{*}_{t},t}\leq\delta$, hence by Eq. (5),
$\big|V_{(N)}(0^{+})-V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\big|<\varepsilon.$
For the first $m_{T}$ rounds we have $|V_{(N)}(0^{+})-V_{(N)}(\cdot)|\leq B$ due to bounded value. Therefore, conditioning on the event in (6),
$\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon.$
Taking expectations and using that the per-round gap is $\leq B$ on the complement of the event in (6), we have
$\mathbb{E}\bigg[\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\bigg]\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon+B\,p_{T}$
$\;\leq\;\frac{O(\sqrt{T}\log\tfrac{1}{\delta})}{T}\,B+\varepsilon+BNTe^{-cO(\sqrt{T}\log\tfrac{1}{\delta})}~\leq~3{\varepsilon}$
for sufficiently large $T$. So we obtain
$\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]\,=\,V_{(N)}(0^{+}).$
Combining with Lemma 6, we conclude that
$V_{(N)}(0^{+})\;\geq\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{I}(\lambda_{1,t}^{I},\ldots,\lambda_{N,t}^{I})\Big]\;=\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{I}\Big],$
which proves Theorem 4.1. ∎
5 Extension to Non-i.i.d. Valuations
We relax the i.i.d. assumption and allow bidder-specific value distributions. In each round $t\in[T]$, bidder $i\in[N]$ draws a value $v_{i,t}\sim F_{i}$ (independently across bidders and rounds), and the outside bid $d^{O}_{t}$ is drawn from $D$. Values and outside bids are bounded by $B>0$. Let $V_{t}^{C}$ and $V_{t}^{I}$ denote the coalition’s total value achieved in round $t$ under coordinated and independent bidding, respectively. Let $U_{i}^{\cdot,A}$ denote the cumulative RoS violation (achieved value minus payment) of bidder $i$ over $T$ rounds under scenario ${\cdot}\in\{C,I\}$ when all bidders follow auto-bidding algorithm $A$.
Total RoS violation.
Our first result shows that, in the non-i.i.d. setting, coordination strictly reduces the total RoS constraint violation, namely, the total utility of the coalition bidders is improved:
Theorem 5.1(Coordination reduces Total RoS Violation, non-iid case)
Assume 1. For any auto-bidding algorithm $A$ that (weakly) overbids, with probability at least $1-\exp\!\big(-\tfrac{\Delta^{2}T}{32B^{2}}\big)$, the total utility $\sum_{i=1}^{N}U_{i}^{C,A}\geq\sum_{i=1}^{N}U_{i}^{I,A}+\frac{\Delta T}{2}$.
See Appendix 0.B for the proof.
Total value.
To compare total value, we require a per-bidder margin condition that ensures each bidder wins with a positive average advantage when they are the highest-value coalition member. For bidder $i$, define $\Delta_{i}:=\mathbb{E}\left[(v_{i}-d^{O})\,\mathbb{I}\{v_{i}=\max_{j\in[N]}v_{j}\}\right]$.
Assumption 2 (Positive per-bidder winning margin)
For every bidder $i\in[N]$, $\Delta_{i}>0$.
Theorem 5.2(Coordination Increases Total Value, non-iid case)
Assume 2, and suppose bidders run Algorithm 3. Then, we have $\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{C}_{t}-\frac{1}{T}\sum_{t=1}^{T}V^{I}_{t}\Big]\geq 0$.
See Appendix 0.C for the proof.
The findings in this section suggest that coordination improves upon independent bidding even in the presence of asymmetry inside the coalition, demonstrating the robustness of our conclusions.
Figure Setting NN $T$ Total Utility (I) Total Utility (C) Total Value (I) Total Value (C)
1(a) i.i.d. 2 4000 -0.011 0.220 0.643 0.666
1(b) i.i.d. 4 4000 -0.077 0.302 0.773 0.800
1(c) i.i.d. 3 4000 -0.048 0.154 0.713 0.749
2(a) non-i.i.d. 2 10000 0.049 0.219 0.715 0.718
2(b) non-i.i.d. 3 20000 -0.014 0.258 0.620 0.633
2(c) non-i.i.d. 5 20000 -0.062 0.619 0.814 0.819
3(a) Real data 4 20000 -0.040 0.156 $\pm$ 0.012 0.614 $\pm$ 0.016 0.929
3(b) Real data 5 20000 -0.065 0.174 $\pm$ 0.012 0.607 $\pm$ 0.012 0.959
Table 1: Performance summary. Confidence intervals below $\pm 0.002$ are omitted.
(a) $N=2,F=U[0,1],D=U[0,0.9]$
(b) $N=4,F=U[0,1],D=U[0,1]$
(c) $N=3,F=U[0,1],D=\mathrm{Beta}(3,2)$
Figure 1: Experiments under i.i.d auto-bidders.
6 Experiments
To complement our theoretical analysis, we conduct a series of experiments evaluating the performance of coordinated auto-bidding, on both synthetic data and a public real-world dataset (iPinYou [ZhangYuanWang2014iPinYou, LiaoEtAl2014iPinYou]).
Unless otherwise stated, all metrics are averaged over 100 independent simulations; shaded regions in figures depict 95% confidence intervals computed as $\bar{x}\pm 1.96\cdot s$, where $\bar{x}$ and $s$ are the sample mean and standard deviation over the 100 simulations.
6.1 Experiment Setup
For all experiments, we run Algorithm 3 with learning rate $\alpha=1/\sqrt{T}$, where $T$ is the total number of auction rounds. The horizon $T$ may vary across experiments because different settings exhibit different convergence rates. The values of $T$ and $N$ (number of coalition bidders) used in each setting are summarized in Table 1.
6.1.1 Synthetic Data
Symmetric setting.
Each coalition bidder $i$ draws values i.i.d. from uniform distribution $U[0,1]$. The outside bid is sampled from $D=U[0,0.9]$ (Figure 1(a)), $D=U[0,1]$ (Figure 1(b)), or $D=\mathrm{Beta}(3,2)$ (Figure 1(c)), representing different market conditions faced by the coalition bidders.
Asymmetric setting.
To stress-test the mechanism under heterogeneity, we sample each bidder’s values from distinct families (Uniform, Beta, or truncated Gaussians with bidder-specific parameters), as summarized in Figure 2. The outside-bid distribution $D$ varies with the scenario: $U[0.2,0.8]$ in Figure 2(a), $\mathrm{Beta}(3,5)$ in Figure 2(b), and $\mathrm{Beta}(2,8)$ in Figure 2(c). Asymmetries arise from both inside the coalition and the outside market conditions.
Full distributional specifications for the synthetic experiments are provided in the figure captions (Figures 1 and 2). All distributions conform to Assumption 1.
6.1.2 Real-World Data
We also experimented with the public iPinYou dataset [ZhangYuanWang2014iPinYou, LiaoEtAl2014iPinYou], released for a real-time bidding competition organized by iPinYou. We use the Season 2 test split, which contains 2,521,6302{,}521{,}630 auction records from $5$ advertisers. Each record provides the bidding (winning) price and the corresponding advertiser ID.
The real-world data impose three limitations: (i) only the winning price is observed, (ii) bidder identities are unavailable, and (iii) there is no label indicating whether any bidders coordinate. To obtain bidder valuations consistent with the data while abstracting away unobserved identities and strategies, we adopt an empirical i.i.d. model: we aggregate all observed winning prices into a common pool and, in each round, draw each coalition bidder’s value independently and uniformly at random with replacement from this pool. This is equivalent to sampling i.i.d. values from the empirical price distribution induced by the dataset.
Following prior work [chen_coordinated_2023], we normalize all prices to $[0,1]$ for numerical stability. The outside bidder’s bid is drawn from the same empirical distribution and then scaled by a random factor from $U[1,2]$ to maintain competitive pressure. We report results for coalitions with $N\in\{4,5\}$ bidders under this setup.
6.2 Results
Metrics.
Across all experiments, we evaluate three metrics for both Independent (Algorithm 1) and Coordination (Algorithm 2): (a) utility (achieved value minus payment), (b) achieved value, and (c) multiplier. Higher utility indicates lower RoS violation; in particular, utility $\geq 0$ implies that the RoS constraint is satisfied. Figures report per-bidder trajectories, whereas Table 1 aggregates to coalition-level totals (sum across bidders). All statistics are averaged over 100 independent runs.
Plot notation.
In the figures, labels of the form bidder $i$ -- I denote bidder $i$ under the Independent scenario, whereas bidder $i$ -- C denotes the Coordination scenario. Shaded regions show 95% confidence intervals over 100 simulations.
Results on synthetic data
In the symmetric settings (Figure 1), coordination sharply increases both each bidder’s utility (hence reducing RoS violation) and the coalition’s total achieved value relative to independent bidding. Although our theory guarantees value improvement at the coalition (total) level, the experiments further show that every individual bidder benefits from coordination by obtaining higher value. Moreover, all multipliers converge to 0, consistent with our theory, and per-bidder statistics (violation, value, and multiplier trajectories) closely track one another due to the i.i.d. value draws.
In the asymmetric settings (Figure 2), the aggregate gains largely persist: coordination improves total value on average and typically improves violations. However, heterogeneity induces dispersion in learning dynamics – bidders with different value distributions exhibit different convergence rates of the multiplier, and individual value improvements are not guaranteed for every bidder even though the coalition-level total value increases. In cases where each bidder has a non-negligible chance to be the winning bidder, we also observe that the multipliers tend to converge toward 0 empirically.
Results on real-world data.
Consistent with the synthetic data experiments, real-world data results show improvements in both individual RoS violation and achieved value, accompanied by multipliers converging steadily toward 0.
7 Discussion
In this paper, we studied coordinated auto-bidding mechanisms for value maximization under RoS constraints, to understand how cooperation among learning agents enhances efficiency in repeated auctions. We proposed a simple yet effective coordination protocol and theoretically proved that coordination improves the coalition’s total value, as well as RoS constraint compliance, under mild assumptions. Our empirical results further validate these findings, showing higher achieved value and consistent convergence patterns. The demonstrated benefit of coordination over independent bidding might provide useful guidance for industrial practice.
Building on these findings, we outline several directions for future work:
  • Auto-bidding outside bidders: While we assumed i.i.d. outside bids (a reasonable assumption when outside bidders are uniformly sampled from a population every round), a natural extension is to incorporate auto-bidding outside bidders, allowing both coalition members and external participants to adjust their bids dynamically.
  • Alternative coordination mechanisms: Another direction is to explore alternative coordination mechanisms, such as highest-bid coordination, to examine how they would affect the bidders’ overall welfare. Highest-bid coordination is analogous to highest-value coordination in the symmetric setting, but not in the asymmetric setting.
  • Other auction formats: Finally, it would be valuable to extend our analysis to other auction formats, especially non-truthful auctions such as first-price and generalized second-price auctions. Such extensions would help test the robustness of our theoretical insights and clarify how coordination interacts with strategic bidding incentives in broader market environments.
Appendix 0.A Proof of Lemma 3
Write the single-bidder’s expected utility as $G(\lambda)=\mathbb{E}_{v\sim F_{(N)}}G(\lambda;v)$ where $G(\lambda;v)$ is the bidder’s expected utility conditioning on having value $v$, which is equal to
$G(\lambda;v)$ $=\mathbb{E}_{d^{O}\sim D}\big[(v-d^{O})\cdot\mathbb{I}[(1+\tfrac{1}{\lambda})v>d^{O}]\big]$
$=vx\big((1+\tfrac{1}{\lambda})v\big)-p\big((1+\tfrac{1}{\lambda})v\big),$
where $x(b)$ is the winning probability when the bidder bids $b$ and $p(b)$ is the expected payment when the bidder bids $b$. Since the second-price auction is a truthful auction, by Myerson’s payment identity for truthful auctions, we have
$p(b)=xb-\int_{0}^{b}x(z)\mathrm{d}z.$ (7)
So,
$G(\lambda;v)$
$=vx\big((1+\tfrac{1}{\lambda})v\big)-(1+\tfrac{1}{\lambda})vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz$
$=-\tfrac{1}{\lambda}vx\big((1+\tfrac{1}{\lambda})v\big)+\int_{0}^{(1+\tfrac{1}{\lambda})v}x(z)dz.$
Take derivative with respect to $\lambda$:
$\tfrac{\partial G(\lambda;v)}{\partial\lambda}$
$=\tfrac{1}{\lambda^{2}}vx\big((1+\tfrac{1}{\lambda})v\big)-\tfrac{1}{\lambda}vx^{\prime}\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})+x\big((1+\tfrac{1}{\lambda})v\big)(-\tfrac{v}{\lambda^{2}})$
$=\tfrac{v^{2}}{\lambda^{3}}x^{\prime}\big((1+\tfrac{1}{\lambda})v\big)$
$\geq 0$
where the last inequality is because $x^{\prime}(b)\geq 0$. Integrating over $v\sim F_{(N)}$, we obtain $G^{\prime}(\lambda)\geq 0$, hence $G(\lambda)$ is increasing.
Appendix 0.B Proof of Theorem 5.1
The argument parallels the proof of Theorem 3.1.
Let $i^{*}_{t}=\arg\max_{i\in[N]}v_{i,t}$ denote the index of the highest-value bidder in round $t$. Conditioned on the history $H_{t-1}$, we analyze the expected increment $\Delta_{i,t}=u_{i,t}^{C,A}-u_{i,t}^{\mathrm{Truth}}$ for bidder $i$:
$\mathbb{E}_{\bm{v}_{t}\sim F,\,d_{t}^{O}\sim D}\!\left[\Delta_{i,t}\mid H_{t-1}\right]$
$=\mathbb{E}_{\bm{v}_{t},\,d_{t}^{O}}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\Delta_{i,t}\mid H_{t-1}\right]$
$=\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\geq d_{t}^{O}]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\right]$
$=\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot(d_{t}^{\mathrm{Truth}}-d_{t}^{O})\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}\leq b_{i,t}^{C,A}]\cdot(v_{i,t}-d_{t}^{O})\right]$
$\geq\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\,\big(\max_{j\neq i}v_{j,t}-d_{t}^{O}\big)_{+}\right]$
$\quad+\mathbb{E}\!\left[\mathbb{I}[i^{*}_{t}=i]\cdot\mathbb{I}[v_{i,t}\leq d_{t}^{O}]\cdot(v_{i,t}-d_{t}^{O})\right].$
Summing over all bidders $i\in[N]$, the indicator $\mathbb{I}[i^{*}_{t}=i]$ ensures that only the highest-value bidder contributes in each round. Hence,
$\mathbb{E}\Big[\sum_{i=1}^{N}\Delta_{i,t}\;\big|\;H_{t-1}\Big]$ $\geq\mathbb{E}\big[(v_{(N-1)}-d_{t}^{O})_{+}\big]-\mathbb{E}\big[(d_{t}^{O}-v_{(N)})_{+}\big]$
$\geq\Delta,$
where the second inequality follows from Assumption 1. Finally, applying Azuma’s inequality (in a similar way as in the proof of Theorem 3.1) completes the proof.
Appendix 0.C Proof of Theorem 5.2
The argument parallels the proof of Theorem 4.1.
Let $G_{i}^{C}(\lambda_{i})$ denote bidder $i$’s expected utility when $v_{i}\sim F_{i}$ and the coalition uses multiplier $\lambda_{i}>0$:
$G_{i}^{C}(\lambda_{i})\;=\;\mathbb{E}_{v_{1}\sim F_{1},\dots,v_{N}\sim F_{N},\;d^{O}\sim D}\big[(v_{i}-d^{O})\,x_{i}(\bm{v},\lambda_{i})\big],$
where $x_{i}(\bm{v},\lambda_{i})=1$ if $v_{i}=\max_{j\in[N]}\{v_{j}\}$ and $(1+\tfrac{1}{\lambda_{i}})v_{i}>d^{O}$, and 0 otherwise.
By reusing the proof of Lemma 3 (Appendix 0.A) with $D$ replaced by the distribution of $\max\{d^{O},\max_{j\neq i}v_{j}\}$, we claim that $G_{i}^{C}(\lambda_{i})$ is increasing in $\lambda_{i}$. Hence,
$G_{i}^{C}(\lambda_{i})\geq\lim_{\mu\to 0^{+}}G_{i}^{C}(\mu)=\mathbb{E}\!\left[(v_{i}-d^{O})\,\mathbb{I}\!\left\{v_{i}=\max_{j\in[N]}\{v_{j}\}\right\}\right]=:\Delta_{i},$
where the limit exists because $G_{i}^{C}(\cdot)$ is bounded and monotone.
We now present the parallel version of Lemma 5.
Lemma 7
Under Assumption 2, for any $i\in[N]$ and $t\geq 1$, with probability at least $1-\exp\!\big(-\tfrac{\Delta_{i}^{2}t}{32B^{2}}\big)$,
$\lambda_{i,t+1}^{C}\;\leq\;\lambda_{i,1}\,\exp\!\Big(-\frac{\Delta_{i}\,t}{2\sqrt{T}}\Big).$
Proof
The expected utility $\mathbb{E}[g_{i,t^{\prime}}]$ is lower bounded by
$\mathbb{E}[g_{i,t^{\prime}}]\;=\;G_{i}^{C}(\lambda_{i,t^{\prime}})\;\geq\;\Delta_{i}.$
Applying Azuma’s inequality (exactly as in the proof of Lemma 5) completes the argument. ∎
Write $v_{(N)}\coloneqq\max_{j\in[N]}v_{j}$ with $v_{j}\sim F_{j}$ mutually independent. Its CDF is $F^{\max}(x)=\Pr[v_{(N)}\leq x]=\prod_{j=1}^{N}F_{j}(x)$. For $\lambda>0$, define
$V_{(N)}(\lambda)$ $\;\coloneqq\;\mathbb{E}_{\,v_{(N)}\sim F^{\max},\,d^{O}\sim D}\Big[\,v_{(N)}\cdot\mathbb{I}\big\{(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big\}\Big],$
$V_{(N)}(0^{+})$ $\;\coloneqq\;\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda).$
The limit $V_{(N)}(0^{+})=\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$ exists because $V_{(N)}(\lambda)$ is monotone and bounded. In the i.i.d. special case $F_{1}=\cdots=F_{N}=F$, $F^{\max}$ coincides with the order-statistic distribution $F_{(N)}$, recovering the original definition.
In the non-i.i.d. case, Lemma 6 still holds:
Lemma 8(copy of Lemma 6)
$V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$.
Proof(Proof of Theorem 5.2)
Fix $\varepsilon>0$ and choose $\delta=\delta(\varepsilon)>0$ as in Equation (5). Let
$\Delta_{\min}\coloneqq\min_{i\in[N]}\Delta_{i}>0,\qquad c_{\min}\coloneqq\frac{\Delta_{\min}^{2}}{32B^{2}}$
Define
$m_{T}\;\coloneqq\;\Big\lceil\frac{2\sqrt{T}}{\Delta_{\min}}\,\log\!\Big(\frac{\lambda_{i,1}}{\delta}\Big)\Big\rceil.$
By Lemma 7, for each $i$ and every $t\geq m_{T}$, $\lambda_{i,t+1}^{C}\leq\delta$ holds with probability at least $1-\exp(-c_{\min}t)$. A union bound over $i=1,\dots,N$ and $t=m_{T},\dots,T$ yields
$\Pr\!\Big[\bigcap_{i=1}^{N}\bigcap_{t=m_{T}}^{T}\{\lambda_{i,t+1}^{C}\leq\delta\}\Big]\;\geq\;1-NT\,e^{-c_{\min}m_{T}}\;\eqqcolon\;1-p_{T}.$
On this event, for all $t\geq m_{T}$ we have $\lambda^{C}_{i^{*}_{t},t}\leq\delta$, hence by (5),
$\big|V_{(N)}(0^{+})-V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\big|<\varepsilon,$
where $i^{*}_{t}=\arg\max_{i\in[N]}v_{i,t}$. For each of the first $m_{T}$ rounds, the absolute gap is $\leq B$. Therefore,
$\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon$
on the above high-probability event; on its complement the gap is $\leq B$. Taking expectations and using $p_{T}\leq NTe^{-c_{\min}m_{T}}$ gives
$\mathbb{E}\Big[\Big|V_{(N)}(0^{+})-\tfrac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\Big]\;\leq\;\frac{m_{T}}{T}B+\varepsilon+B\,p_{T}\;\xrightarrow[T\to\infty]{}\;0.$
Hence
$\lim_{T\to\infty}\mathbb{E}\Big[\tfrac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big]=V_{(N)}(0^{+}).$
Finally, combining with Lemma 8 yields the desired inequality against the independent-bidding benchmark, completing the proof. ∎
Appendix 0.D Additional Experiment Results
This appendix presents additional figures in the experiments: Figure 2 for the asymmetric setting on synthetic data; Figure 3 for the real-world dataset.
(a) $F_{1}=\mathrm{Beta}(2,5),F_{2}=\mathrm{Beta}(5,2),D=U[0.2,0.8],N=2$
(b) $F_{1}=\mathrm{Beta}(2,6),F_{2}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.6,0.15),F_{3}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.4,0.2),D=\mathrm{Beta}(3,5),N=3$
(c) $F_{1}=U[0.2,0.8],F_{2}=\mathrm{Beta}(4,3),F_{3}=\mathrm{Beta}(6,2),F_{4}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.5,0.1),F_{5}=\mathcal{N}_{\mathrm{cap}=[0,1]}(0.7,0.12),D=\mathrm{Beta}(2,8),N=5$
Figure 2: Experiments under non-i.i.d auto-bidders.
(a) $N=4$
(b) $N=5$
Figure 3: Experiments in real-world datasets.